# HDO ice detected toward an isolated low-mass protostar with JWST

### v2

@@ 1 Introduction
Interferometric (sub)millimeter gas-phase studies thus attempted to quantify interstellar ice HDO/H2O ratios indirectly instead by targeting the warm (100 K) inner regions of protostellar envelopes (“hot corinos”). Because it is often argued that the majority of the gas species detected in these environments originate from the thermal desorption of ices with little subsequent gas-phase chemical alteration (e.g., Requena-Torres et al. 2006; Coletta et al. 2020; Van Gelder et al. 2020; Nazari et al. 2021; Chen et al. 2023), their measured gas HDO/H2O ratios are assumed by some to reflect the ratios of their ice envelopes. Toward four protostars located in the clustered cloud complexes NGC 1333 and Ophiuchus, the measured gas HDO/H2O ratios range between 5-910-4 (Persson et al., 2014), very similar to the ratios measured in several comets in our solar system (e.g., Eberhardt et al. 1995; Villanueva et al. 2008; Altwegg et al. 2015). Because our solar system likely formed in a clustered star-forming region (Adams, 2010; Desch & Miret-Roig, 2024), this similarity has been used to assert the claim that cometary water ice was directly inherited from the prestellar cloud or protostellar envelope with minimal chemical processing in the protoplanetary disk in which comets formed (Cleeves et al., 2014). In contrast, the gas HDO/H2O ratios measured toward five distributed hot corinos found in isolated environments (i.e., they are not associated with star-forming cloud complexes) tend to be a factor of 2-10 higher (Jensen et al., 2019, 2021). These studies suggested that such a difference in gas HDO/H2O ratios could be due to the lower radiation experienced by more isolated star-forming regions, which implies colder and/or longer prestellar stages than those in more clustered star-forming regions. It must be emphasized, however, that this reported dependence of the protostellar gas HDO/H2O on the local environment is based on a very limited sample size of sources, as discussed further in Section 3.3.
@@
The advent of JWST, with its unprecedented sensitivity, offers an opportunity to revisit the search for HDO ice toward such star-forming regions. Recently, Slavicinska et al. (2024) demonstrated the capability of JWST to unambiguously detect HDO ice toward two massive protostars and subsequently calculated the ice HDO/H2O ratios of their envelopes. Here, we focus on the source L1527 IRS (also IRAS 04368+2557, hereafter L1527), an isolated low-mass class 0/I protostar (Tobin et al., 2013) found in the Taurus molecular cloud ( = 140 pc) that is deeply embedded within a dense 0.9 M⊙ envelope and has a 0.0075 M⊙, 75-125 AU-sized disk observed in the edge-on configuration (Aso et al., 2017; Kristensen et al., 2012; Tobin et al., 2013). The protostar currently has a mass of 0.2 M⊙ but may eventually accrete enough material from its envelope to grow to a sun-like mass (Tobin et al., 2012).
@@ 5 Acknowedgments
K.S. thanks Brian Ferrari and Gary Huss for helpful scientific discussions. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs #1798, #1290, and #1960. All the data used in this paper can be found in MAST: http://dx.doi.org/10.17909/jhkb-w045 (catalog 10.17909/jhkb-w045). The following National and International Funding Agencies funded and supported the MIRI development: NASA; ESA; Belgian Science Policy Office (BELSPO); Centre Nationale d’Etudes Spatiales (CNES); Danish National Space Centre; Deutsches Zentrum fur Luft- und Raumfahrt (DLR); Enterprise Ireland; Ministerio De Economiá y Competividad; Netherlands Research School for Astronomy (NOVA); Netherlands Organisation for Scientific Research (NWO); Science and Technology Facilities Council; Swiss Space Office; Swedish National Space Agency; and UK Space Agency. Astrochemistry at Leiden is supported by funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101019751 MOLDISK), the Netherlands Research School for Astronomy (NOVA), and the Danish National Research Foundation through the Center of Excellence “InterCat” (Grant agreement no.: DNRF150). A.C.G. acknowledges support from PRIN-MUR 2022 20228JPA3A “The path to star and planet formation in the JWST era (PATH)” funded by NextGeneration EU and by INAF-GoG 2022 “NIR-dark Accretion Outbursts in Massive Young stellar objects (NAOMY)” and Large Grant INAF 2022 “YSOs Outflows, Disks and Accretion: towards a global framework for the evolution of planet forming systems (YODA)”.
@@ B.2 Global continuum and silicate subtraction
To extract the full combined NIRSpec+MIRI spectrum into the optical depth scale, we defined two separate but connecting global continua in the near-IR (2.5-5 m) and mid-IR (5-28 m) ranges by fitting low-order polynomials to the data on a log(Fλ) versus log() scale, similar to the procedure described in Boogert et al. (2008). The resulting global continuum fit to the 2.5-27.9 m spectral range is shown in Figure B.2. Similarly curved SEDs have been used to subtract the global continuum from other protostellar spectra (e.g., see the CrA IRAS 32, IRAS 03235+3004, and CRBR 2422.8-3423 global continua in Figure 1 of Boogert et al. 2008) and can be reproduced by radiative transfer models of protostellar systems (Crapsi et al., 2008). In the case of our global continuum for L1527, the polynomial continuum fit can also be reproduced well with two blackbody curves, one at 450 K and another at 100 K, roughly simulating dust emission in two different regions of the warm L1527 envelope.
@@ Appendix D Ice fitting uncertainties
Although SH-bearing ice species are not the focus of this work, we report in Section 3.1 H2S and NH4SH upper limits (0.3 and 13% w.r.t. H2O, respectively) calculated from the Gaussian fit to the excess at 3.9 m, in the case that these upper limits may be useful for other studies of sulfur in star-forming regions or interstellar ices. However, it is evident from Figure D.1 that the optical depth of this Gaussian can vary dramatically with the continuum choice, resulting in high uncertainties in these reported upper limits. Depending on the continuum, the H2S upper limits can range from 0.07-2%. For NH4SH, the upper limit is 4% in the case of the lower limit continuum choice shown in the bottom left panel of Figure D.1. In the case of the upper limit continuum choice in the bottom right panel, the 3.9 m Gaussian profile is too broad to be attributed exclusively with NH4SH. Furthermore, the L1527 NH4+ column density derived from the 6.85 m feature provides an additional constraint that the NH4SH upper limit cannot exceed that derived from the fit presented in Figure 2 (13%).

### v3

@@ 1 Introduction
Interferometric (sub)millimeter gas-phase studies thus attempted to quantify interstellar ice HDO/H2O ratios indirectly instead by targeting the warm (100 K) inner regions of protostellar envelopes (“hot corinos”). Because it is often argued that the majority of the gas species detected in these environments originate from the thermal desorption of ices with little subsequent gas-phase chemical alteration (e.g., Requena-Torres et al. 2006; Coletta et al. 2020; Van Gelder et al. 2020; Nazari et al. 2021; Chen et al. 2023), their measured gas HDO/H2O ratios are assumed by some to reflect the ratios of their ice envelopes. Toward four protostars located in the clustered cloud complexes NGC 1333 and Ophiuchus, the measured gas HDO/H2O ratios range between 5-910-4 (Persson et al., 2014), very similar to the ratios measured in several comets in our solar system (e.g., Eberhardt et al. 1995; Villanueva et al. 2008; Altwegg et al. 2015). Because our solar system likely formed in a clustered star-forming region (Adams, 2010; Desch & Miret-Roig, 2024), this similarity has been used to assert the claim that cometary water ice was directly inherited from the prestellar cloud or protostellar envelope with minimal chemical processing in the protoplanetary disk in which comets formed (Cleeves et al., 2014). In contrast, the gas HDO/H2O ratios measured toward five distributed hot corinos found in isolated environments (i.e., they are not associated with star-forming cloud complexes) tend to be a factor of 2-10 higher (Jensen et al., 2019, 2021; Andreu et al., 2023; Tobin et al., 2023). Some of these studies suggested that such a difference in gas HDO/H2O ratios could be due to the lower radiation experienced by more isolated star-forming regions, which implies colder and/or longer prestellar stages than those in more clustered star-forming regions. It must be emphasized, however, that this reported dependence of the protostellar gas HDO/H2O on the local environment is based on a very limited sample size of sources, as discussed further in Section 3.3.
@@
The advent of JWST, with its unprecedented sensitivity, offers an opportunity to revisit the search for HDO ice toward such star-forming regions. Recently, Slavicinska et al. (2024) demonstrated the capability of JWST to unambiguously detect HDO ice toward two massive protostars and subsequently calculated the ice HDO/H2O ratios of their envelopes. Here, we focus on the source L1527 IRS (also IRAS 04368+2557, hereafter L1527), an isolated low-mass class 0/I protostar (Tobin et al., 2013) found in the Taurus molecular cloud ( = 140 pc) that is deeply embedded within a dense 0.9 M⊙ envelope and has a 0.0075 M⊙, 75-125 AU-sized disk observed in the edge-on configuration (Aso et al., 2017; Kristensen et al., 2012; Tobin et al., 2013). The protostar currently has a mass around 0.3-0.5 M⊙ (Aso et al., 2017; van’t Hoff et al., 2023) but may eventually accrete enough material from its envelope to grow to a sun-like mass (Tobin et al., 2012).
@@ 5 Acknowedgments
K.S. thanks Brian Ferrari and Gary Huss for helpful scientific discussions. The authors thank an anonymous reviewer for their insightful comments that helped to improve this paper. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs #1798, #1290, and #1960. All the data used in this paper can be found in MAST: http://dx.doi.org/10.17909/jhkb-w045 (catalog 10.17909/jhkb-w045). The following National and International Funding Agencies funded and supported the MIRI development: NASA; ESA; Belgian Science Policy Office (BELSPO); Centre Nationale d’Etudes Spatiales (CNES); Danish National Space Centre; Deutsches Zentrum fur Luft- und Raumfahrt (DLR); Enterprise Ireland; Ministerio De Economiá y Competividad; Netherlands Research School for Astronomy (NOVA); Netherlands Organisation for Scientific Research (NWO); Science and Technology Facilities Council; Swiss Space Office; Swedish National Space Agency; and UK Space Agency. Astrochemistry at Leiden is supported by funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101019751 MOLDISK), the Netherlands Research School for Astronomy (NOVA), and the Danish National Research Foundation through the Center of Excellence “InterCat” (Grant agreement no.: DNRF150). A.C.G. acknowledges support from PRIN-MUR 2022 20228JPA3A “The path to star and planet formation in the JWST era (PATH)” funded by NextGeneration EU and by INAF-GoG 2022 “NIR-dark Accretion Outbursts in Massive Young stellar objects (NAOMY)” and Large Grant INAF 2022 “YSOs Outflows, Disks and Accretion: towards a global framework for the evolution of planet forming systems (YODA)”.
@@ B.2 Global continuum and silicate subtraction
To extract the full combined NIRSpec+MIRI spectrum into the optical depth scale, we defined two separate but connecting global continua in the near-IR (2.5-5 m) and mid-IR (5-28 m) ranges by fitting low-order polynomials to the data on a log(Fλ) versus log() scale, similar to the procedure described in Boogert et al. (2008). The resulting global continuum fit to the 2.5-27.9 m spectral range is shown in Figure B.2. Similarly curved SEDs have been used to subtract the global continuum from other protostellar spectra (e.g., see the CrA IRAS 32, IRAS 03235+3004, and CRBR 2422.8-3423 global continua in Figure 1 of Boogert et al. 2008) and can be reproduced by radiative transfer models of protostellar systems (Crapsi et al., 2008). In the case of our global continuum for L1527, the polynomial continuum fit can also be reproduced well with two blackbody curves, one at 450 K and another at 100 K, roughly simulating scattered light and thermal emission from different regions of the L1527 system.
@@ Appendix D Ice fitting uncertainties
Although SH-bearing ice species are not the focus of this work, we report in Section 3.1 H2S and NH4SH upper limits (0.3 and 13% w.r.t. H2O, respectively) calculated from the Gaussian fit to the excess at 3.9 m, in the case that these upper limits may be useful for other studies of sulfur in star-forming regions or interstellar ices. However, it is evident from Figure D.1 that the optical depth of this Gaussian can vary dramatically with the continuum choice, resulting in high uncertainties in these reported upper limits. Depending on the continuum, the H2S upper limits can range from 0.07-2%. For NH4SH, the upper limit is 4% in the case of the lower limit continuum choice shown in the bottom left panel of Figure D.1. In the case of the upper limit continuum choice in the bottom right panel, the 3.9 m Gaussian profile is too broad to be attributed exclusively with NH4SH. Furthermore, the L1527 NH4+ column density derived from the 6.85 m feature (Slavicinska et al., 2025) provides an additional constraint that the NH4SH upper limit cannot exceed that derived from the fit presented in Figure 2 (13%).